A toy rocket is launched from ground level. Its distance in feet from the ground in t seconds is given by
s (t) = - 16 t 2 + 208 t
At what times will the rocket be 550 feet from the ground?

Answer by Alan3354(69443)

A toy rocket is launched from ground level. Its distance in feet from the ground in t seconds is given by
s (t) = - 16 t 2 + 208 t
At what times will the rocket be 550 feet from the ground?
------------
s(t) = -16t^2 + 208 t
s(t) = -16t^2 + 208 t = 55
-16t^2 + 208 t - 55 = 0
Solve for t
It'll be at 550 ft going up, and again coming down.